# 3D Anderson localization of light in disordered systems of dielectric particles

### v1

@@ Abstract
We present the results of full wave numerical simulations of light transmission by layers of irregular dielectric particles that demonstrate 3D Anderson localization of light in disordered uncorrelated discrete media. Our simulations show that a high degree of disorder in a dense layer prevents transverse spreading of a propagating beam and allows anomalous conservation of its initial polarization state. A transition from the purely diffusive regime to a non-exponential time dependence when reaching the Ioffe-Regel criterion is observed in short pulse transmission measurements. The effect depends on the turbidity of the layer: increasing the volume fraction and the refractive index leads to the enhancement of the non-exponential behaviour caused by the disorder in the medium which is a clear signature of Anderson localization.
keywords:
disordered medium, light localization, light scattering
1 Introduction
Complete understanding and control of light propagation in disordered media can be advantageous in different applications in photonics like imaging and focusing in random media, random lasing, radiative transfer in particulate media and optical remote sensing [1, 2, 3, 4, 5]. However, this goal has not been achieved yet, in particular, for dense turbid media, due to the complexity of the problem for both analytical analysis and numerical simulations.
Light propagation in sparse media, in which scatterers are separated by distances larger than the wavelength, can be understood as a sequence of scattering events that are determined by the optical properties of individual particles, like single-scattering phase function and single-scattering albedo. If the number of such events is large enough, the optical transport process can be modelled using analytical multiple-scattering theories and the classic diffusion equation (e.g., [6, 7, 8]). In dense media with high concentration of scatterers, multiple scattering includes near-field interactions, and more complex transport phenomena may take place. For example, interference of counter-propagating waves may result in a weak localization and a coherent backscattering enhancement [9, 10]. Strong Anderson localization (AL) has been predicted for media with strong disorder and with multiple scattering not being suppressed by absorption [11, 12]. An important condition for the AL emergence is the sub-wavelength scale of coherent multiple scattering that can be expressed through the Ioffe-Regel criterion $kl^{*}<1$ for the mean free path length $l^{*}$ (where $k$ is the wavenumber) [13, 14]. This suggests a necessity of dense packing of particles in a disordered system in order to obtain AL. With high enough scattering strength of dielectric scatterers, that can be controlled by the variation of their refractive index, one can expect signatures of AL in the observations of light transmission by layers of particles. However, even non-absorbing layers with required properties and thickness $kL\gg 1$ can be quite opaque as it was already recognized by Anderson [11], which complicates the study of their transmission properties even further.
Generally, AL is characterized by a transition from diffusive propagation to a reduction and even complete halt of diffusion. An explanation involving trajectories with closed loops and increased return probability is often suggested [15, 16, 17]. However, the actual mechanism behind such dynamics has not been described yet in detail. The phenomenon is general in nature and can be observed for different kinds of waves, e.g., in acoustics [18], quantum systems [19, 20] and 2D photonic structures [21]. At the same time, the existence of 3D AL for electromagnetic waves in disordered uncorrelated media remains under debate. Experimental measurements of light transmission by layers of highly scattering particulate materials could shed light on the problem [17, 22, 23]. However, certain problems with precise control of the sample properties and accounting for the effects of absorption [24] and inelastic scattering [25, 26, 27] make this difficult. The most recent experimental results reported in [23] are debated because of ambiguous interpretation of the detected signature of AL that can be explained by fluorescence [27].
In theoretical studies of AL, one has to properly describe the disordered structure of a model medium and carefully take into account the multiple scattering near-field effects, including the polarization state of the scattered near field [28, 29, 16, 30]. Conventional multiple scattering theories cannot be applied here because of the small mean free path values at dense packing where the far-field single scattering properties of particles do not play role. In existing analytical models of the AL phenomenon, simplified descriptions of the media are used, e.g., a cloud of point scatterers [31] or a dielectric constant fluctuation around the average value $\epsilon(\vec{r})=\bar{\epsilon}+\delta\epsilon(\vec{r})$ [15]. If applied to the real ’white paint’-like 3D media, this can be insufficient.
Full wave numerical simulations can avoid approximations and unnecessary scattering effects considering pure interaction of the electromagnetic field with the target structure that may have arbitrary geometry. Modern numerical methods allow generation of particulate layer samples with high degrees of disorder, arbitrary shape of scatterers and packing densities [28, 32, 33]. At the same time, such an electromagnetic problem is multi-scale by definition. If the target object is much larger than the wavelength and has sub-wavelength structural features, accurate solutions require large computational meshes. This is a technical problem that can be resolved with the application of high-performance computing (HPC). However, only a few works on 3D numerical simulations and light localization in disordered and partially disordered structures have been reported [34, 35, 36, 37].
@@ 1 Introduction
Observable characteristics of AL are the reduction of transverse spreading of the propagating light beam [38, 36, 34] and, in particular, time-delayed transmitted energy decay [39, 34]. These anomalies as unambiguous signatures of AL were found for the first time in a recent full wave numerical study of light transmission through 3D layers of densely packed perfect electric conductor (PEC) spheres [34]. That work suggests a possibility of the AL existence for electromagnetic waves in 3D in general. However, the case of dielectric non-absorbing spheres in layers with similar parameters and refractive indices up to $n=10$ showed only diffuse transport. Thus, the question whether AL can be observed for electromagnetic waves for disordered dielectric 3D structures remains an open problem to date.
@@
In this work, we apply numerical simulations and HPC to study if conditions for AL can be created at $kl^{*}\lesssim 1$, according to the Ioffe-Regel criterion, and if AL signatures can be observed in light transmission by layers of dielectric particles. Reaching high enough degree of disorder at dense packing of irregular constituents, we observe a transition of classical diffusion characterised by the exponential time dependence of transmission $T(t)$ to AL with the $T(t)$ decay rate deviating from exponent. This behaviour is accompanied by the reduction of transverse spreading observed in the propagation of a focused beam. Thus, we demonstrate a possibility of AL in the ’white-paint’-like 3D systems of dielectric particles.
2 Numerical model
@@
We apply the discontinuous Galerkin time-domain method (DGTD) [40, 32, 41, 42] to solve Maxwell’s equations for multi-particle layers illuminated by a plane wave as well as by a focused beam. For a qualitative benchmark of the DGTD light transmission computations for simplified structures we utilized the CST Microwave Studio software package, which is based on the Time Domain Finite Integration Technique (FIT) (see Supplementary material A.2).
@@
We note that the $kl^{*}\lesssim 1$ criterion unconditionally requires dense packing of scatterers. Maximally random jammed packings of regular constituents, like spheres and regular polyhedra, always form hyper-uniform structures regardless of the packing algorithm [43, 44]. This kind of discrete media packed near their corresponding maxima, that depend on the type of constituents, have a reduced degree of disorder which influences their transmission properties (e.g., [45, 28] ). The problem can be solved by considering scatterers with random irregular shapes. To our knowledge, numerical simulations studying light localization in such systems have not been reported yet. Observation of AL in 3D layers of PEC spheres and its absence in the dielectric equivalent structures [34] provides us with an additional motivation to consider the case of irregular constituents.
@@
The constituent particle size is determined by the goal of satisfying the $kl^{*}\lesssim 1$ criterion and, possibly, to create conditions for light localization. The scale of free-space voids in a mono-disperse system becomes comparable to the particle size at large volume fractions. We take the particle size $kr=1.1$, where $r$ is the radius of the circumscribing sphere. In densely packed samples of compact irregular shapes their circumscribing spheres overlap making the distances between neighbour particles sufficiently small.
We consider two types of cylindric layers: thick ones for studying the focused beam propagation and thinner, but horizontally more extended ones for time-resolved transmission measurements. In the first case, the diameter and thickness are $kR=22.1$ and $kL=19.5$, respectively, and the number of particles is 20000 (Fig. 1a). In the second case, layers with various packing densities $\rho$ have larger diameter to thickness ratios and numbers of particles up to 19000. A layer sample with $kR=32$, $kL=14.5$ and $\rho=0.44$ and its vertical cross-section are shown in (Figs. 3ab). We note that the sample sizes are limited by our computational resources and the complexity of the unstructured mesh generation. For the DGTD simulations, we generate meshes with the particle systems surrounded by the perfectly matching layer (PML) to simulate an open boundary. We do not apply periodic boundary conditions in order to avoid possible uncertainties caused by the periodicity of the mesh. Cylindrical symmetry of samples may also affect the result. Therefore, we use samples with two types of the side boundaries: a randomly perturbed one and a smooth boundary. In the first case defects of the scale of $\sim\lambda$ that break such symmetry are introduced. In the second case a smooth boundary forms a dielectric discontinuity between the disordered medium and free space and, generally, one can expect whispering-gallery (WG) modes due to total internal reflection and cylindrical symmetry [48].
@@ 3 Results and discussion
Figure 1: Propagation of a continuous, focused beam through a dense and a sparse layers of irregular particles. (a) and (c), samples with volume fractions of 0.48 (dense) and 0.16 (sparse), respectively. (b) and (d), steady-state near-field intensity distributions $|E|^{2}$ in XY, YZ and back-XZ planes for both layers. A white dotted line indicates a layer boundary with the diameter $kR=22.1$ and the thickness $kL=19.5$. The incident beam is $E_{x}$-polarized. The size of particles is $kr=1.1$, and the refractive index is $n=3.0$. The sparse structure shows a classic multiple-scattering pattern while dense packing strongly limits transverse spreading of the beam. An animation of a short pulse propagation, illustrating the transport mechanism in the dense case, is provided in the Supplementary material A.1.
@@
Diffusive propagation of light is characterized by a transverse spread of energy and exponential energy decay in time as a result of multiple scattering. The geometry of the field spread experiences a change when particles are densely packed. This effect depends on the particle size and refractive index. Generally, at dense packing, far-field single scattering becomes impossible, and classic multiple scattering turns into percolation of light along random transmission channels [28, 29, 30]. Depending on the particle size and degree of disorder, this can lead to formation of a Caylee-tree-like channel [28] or a transversely localized propagation [38, 36]. This may occur even at a relatively small refractive index $n=1.5$ and $l^{*}\lesssim\lambda$ [28], which, however, should be insufficient for AL. Nevertheless, the diffuse field component appears to be significantly reduced under such conditions, which can be seen in the polarimetric measurements of backscattering from high-albedo powders with different packing densities [49, 50, 51]. Little free space between the neighboring scatterers in a medium reduces the role of diffraction after each scattering event and, consequently, the contribution of the diffuse field [28]. The measured polarimetric features produced by single particles on the surface, that are suppressed by unpolarized diffuse multiple scattering in sparse samples, get enhanced in compressed ones. Hence, in the search for real AL, we consider samples with parameters that are close to those predicted in the theoretical works.
@@
In Fig. 1, we study focused beam propagation through a thick layer of particles with a size $kr=1.1$ and a refractive index $n=3.0$. The internal field patterns are compared for a dense and a sparse structures of $\rho=0.48$ (Fig. 1a) and 0.16 (Fig. 1c), respectively. A continuous $E_{x}$-polarized focused beam incident on the front side in each case propagates in $Z$ direction. Steady-state intensity distributions as cross-sections in different planes are shown in Figs. 1b and 1d. The ”XY” panels correspond to the cross-sections near the back sides of the samples.
@@
In the sparse case (Fig. 1d), we observe classic scattering by particle groups and individual particles, which appear to be weak scatterers at large separations from each other. In a dense system, the size of voids is reduced to $kl^{*}\lesssim 1$, and a focused beam tends to preserve its diameter along the path, re-emitting transmitted energy in the center of the back side. Small isotropic spreading occurs due to the leakage of energy through the evanescent field coupling. Therefore, the intensity of the beam decreases, but its geometric cross-section remains constant at the distance of the layer thickness. The dynamics of the process can be seen in the animation in the Supplementary material A.1. It shows the evolution of a focused beam, percolating through a dense layer without transverse spreading. Light propagation evolves mainly through evanescent coupling between individual scatterers without conventional free-space transport of the scattered waves. This reminds us of an evolution of a cellular automata (CA) in which the vector field state in an elementary cell at each time step is determined by its nearest environment, i.e. the shape of neighbour particles or the shape of a sub-wavelength void that is formed by the neighbors. The laws of interaction of the vector field and the air-dielectric interface become the CA rules. Then, a sequence of such interactions determines the evolution of the entire vector field from the initial state, resulting in a forward pulse propagation.
The fundamental property of the observed transport mechanism is that most of the energy of the propagating beam is concentrated in the free-space cavities, and little is refracted into the dielectric particles. A wave can couple to the neighbor-particle interface while being still coupled to the previous one propagating in a void between two interfaces. On the other hand, these interfaces are formally disconnected and belong to different scatterers. Thus, the field appears to be constantly coupled and one can say that there are no more ”scattering events”. Therefore, the term ”free path length” becomes irrelevant in its classic definition (see also [14]). $kl^{*}<1$ implies also an extremely small mean free time $t^{*}$, and the near field, being coupled to a particle, most often reaches its neighbour at time $t\ll T$, where $T$ is the wave oscillation period. Thus, we refer to the $kl^{*}$ parameter only to characterise the size of the voids in a particulate medium with respect to the wavelength.
Furthermore, we note that the wave packet is partially preserved along the propagation path. Its effective wavelength becomes smaller, $\lambda_{eff}\approx 0.6\lambda_{inc}$, increasing the effective size of particles to $k_{eff}r\approx 1.8$. Nevertheless, the condition $kl^{*}\lesssim 1$ is preserved for most of the voids.
Figure 2: Propagation of a continuous pulse through the dense layer shown in Fig. 1a. Steady state distribution of the components of the electric near field $\vec{E}$ (XZ-plane) show that the transmitted pulse is not totally depolarized but, on the contrary, preserves most of its initial polarization state $E_{x}$ after transmission. The $E_{y}$ component as well as $E_{z}$ is excited locally in the voids between the particles along the propagation path but is not transmitted directionally.
Diffusive propagation in a disordered medium depolarizes incident radiation after many scattering events that are strongly stochastic. A steady-state distribution of the electric field component $E_{x}$ in Fig. 2 computed for a dense layer shows a counter-intuitive conservation of the original polarization state of the propagating beam at the exit. The $E_{y}$ and $E_{z}$ components are excited along the path and play a role in the field coupling process but they do not propagate directionally. This fact can explain the halt of transverse spreading: the propagating beam ”prefers” cavities where the initial polarization direction can be preserved and, hence, the $\vec{k}$-vector bending is avoided and the direction of propagation is not changed.
@@
Figure 3: (a) A layer of 19000 irregular particles packed with a volume fraction of 0.44 and with dimensions of $kR=32$ and $kL=14.5$. The size of particles is $kr=1.1$ and the refractive index is $n=3.0$. (b) vertical cross-section of the layer, showing intentional defects on the side boundary and the characteristic scale of voids $kl\lesssim 1$. (c) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.17$ to $0.5$ as a function of time $T(t)$, exhibiting a transition from the exponential regime to a non-exponential decay at longer times. (d) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. Both axes have logarithmic scale. A dense layer with $\rho=0.44$ demonstrates a possible transition to $t^{-1}$ (dashed line) dependence at times $t>20t_{D}$.
Figure 4: Simulation results for layers with different volume fractions and smooth side boundaries. The sample dimensions are of $kR=26$ and $kL=12$, respectively. The size of particles is $kr=1$ and the refractive index is $n=3.0$. (a) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.1$ to 0.5 as a function of time $T(t)$, exhibiting a transition from the purely exponential regime to a non-exponential decay at $\rho\gtrsim 0.3$. (b) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. Dense layers demonstrate a $t^{-1}$ (dashed lines) dependence at times $t>20t_{D}$. (c) and (d), Snapshots of the internal electric field (XY-cross-sections, $E_{x}$ component) distributions showing a whispering-gallery wave propagating along the boundary at time $t=25t_{D}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers. In the last case $T(t)$ is exponential despite the whispering-gallery effect.

### v4

@@ Abstract
We present the results of full-wave numerical simulations of light transmission through layers of irregular dielectric particles, demonstrating three-dimensional Anderson localization of light in disordered, uncorrelated discrete media. Our simulations show that a high degree of disorder in a dense layer suppresses the transverse spreading of a propagating beam. A transition from the purely diffusive regime to a non-exponential temporal dependence is observed in short-pulse time-resolved transmission measurements as the system approaches the Ioffe-Regel condition. Along with this, near-field dynamics leads over time to the formation of spatially localized modes and the transmission spectrum becomes consistent with the Thouless criterion. The effect depends on the turbidity of the layer: increasing the volume fraction of scatterers and the refractive index contrast enhances the non-exponential behavior induced by disorder, which is a clear signature of Anderson localization.
preprint: APS/123-QED
A complete understanding and control of light propagation in discrete disordered media can be advantageous in different applications in photonics, like imaging and focusing in random media, random lasing, radiative transfer and optical remote sensing [1, 2, 3, 4, 5]. However, modeling light transport phenomena in highly disordered turbid media remains a challenge due to the inherent complexity and multi-scale nature of the problem. While classic wave diffusuion and weak localization, observed as coherent backscattering enhancement [6, 7, 8], are well understood, strong Anderson localization of light (AL) in 3D structures is a subject of ongoing research. AL has been predicted for media with strong disorder and strong multiple scattering [9, 10]. Generally, it is characterized by a transition from diffusive propagation to a reduction and even complete halt of diffusion. An explanation involving trajectories with closed loops and increased return probability is often suggested [11, 12, 13]. However, the actual mechanism behind such dynamics has not been described yet in detail. An important condition for the AL emergence is the sub-wavelength scale of coherent multiple scattering that can be expressed through the Ioffe-Regel criterion $kl^{*}\lesssim 1$ for the mean free path length $l^{*}$ (where $k$ is the wavenumber) [14, 15]. The phenomenon is general in nature and can be observed for different kinds of waves, e.g., in acoustics [16], quantum systems [17, 18], 2D [19, 20] and spatially correlated 3D photonic structures [21]. The existence of AL for electromagnetic waves in 3D uncorrelated media and metallic materials has been demonstrated recently through a numerical study [22].
@@
Observable characteristics of AL are the reduction of transverse spreading of the propagating light beam [23, 24, 22] and, in particular, time-delayed transmitted energy decay [25, 22]. Transmission spectrum in this case is characterized by non-overlapping sharp peaks [26, 27, 28, 29]. Numerical modeling of light transmission through layers of overlapping metallic spheres revealed these anomalies as unambiguous signatures of AL [22], suggesting its general possibility in 3D. At the same time analogous dielectric structures did not exhibit the phenomenon [22]. Thus, the question whether AL can be observed for electromagnetic waves and disordered dielectric 3D media remains an open problem to date. Experimental measurements of light transmission by layers of highly scattering particulate materials could shed light on the problem [13, 30, 31]. However, certain problems with precise control of the sample properties and accounting for the effects of absorption [32] and inelastic scattering [33, 34, 35] make this difficult.
@@
In this work, we apply numerical simulations and high-performance computing to demonstrate that AL signatures can be observed in light transmission by 3D layers of irregular dielectric particles. Reaching high enough degree of spatial disorder, we obtain a transition of classical diffusion characterized by the exponential time dependence of transmission $T(t)$ to AL with the $T(t)$ decay rate deviating from exponent. This behavior is accompanied by the reduction of transverse spreading observed in the propagation of a focused beam.
We note that the problem requires proper description of the disordered geometry of a model medium and carefull accounting for the multiple scattering near-field effects, including the polarization state of the scattered near field [36, 37, 12, 38, 39]. In existing analytical solutions, simplified descriptions of the media are used (e.g., [39, 11, 40]) which can be insufficient if applied to the vector electromagnetic field in real ’white paint’-like 3D media. With full wave numerical simulations one can avoid approximations and consider pure electromagnetic scattering in the target structure with arbitrary geometry [36, 41, 42, 22, 43, 24, 21].
@@
We use the discontinuous Galerkin time-domain method (DGTD) [44, 41, 45, 46] to solve Maxwell’s equations for multi-particle layers illuminated by a plane wave as well as by a focused beam (see details in Supplementary material I). For a qualitative benchmark of the DGTD computations for simplified structures we utilized the CST Microwave Studio software package, which is based on the Time Domain Finite Integration Technique (FIT) (see Supplementary material II).
@@
The $kl^{*}\lesssim 1$ criterion unconditionally requires dense packing of scatterers. Maximally random jammed packings of regular constituents, like spheres and regular polyhedra, always form hyper-uniform structures regardless of the packing algorithm [47]. This kind of discrete media packed near their corresponding maxima have a reduced degree of disorder which influences their transmission properties (e.g., [48, 36] ). The problem can be solved by considering scatterers with random irregular shapes.
@@
The constituent particle size is determined by the goal of satisfying the $kl^{*}\lesssim 1$ criterion. The scale of free-space voids in a mono-disperse system becomes comparable to the particle size at large volume fractions. We take the particle size $X_{r}=kr=1.1$, where $r$ is the radius of the circumscribing sphere. In densely packed samples of compact irregular shapes their circumscribing spheres overlap making the distances between neighbour particles sufficiently small.
We consider two types of cylindric layers: thick ones for studying the focused beam propagation and thinner, but horizontally more extended ones for time-resolved transmission measurements. In the first case, the dimensionless size parameters of the diameter and thickness are $X_{R}=22.1$ and $X_{L}=19.5$, respectively, and the number of particles is 20000 (Fig. 1a). In the second case, layers with various volume fractions $\rho$ have larger diameter to thickness ratios and numbers of particles up to 19000. A layer sample with $X_{R}=32$, $X_{L}=14.5$ and $\rho=0.44$ and its vertical cross-section are shown in Figs. 2ab. For the DGTD simulations, we generate meshes with the particle systems surrounded by the perfectly matching layer (PML) to simulate an open boundary. We do not apply periodic boundary conditions in order to avoid possible uncertainties caused by the periodicity of the mesh. Cylindrical symmetry of samples may also affect the result. Therefore, we use samples with two types of the side boundaries: a randomly perturbed one and a smooth boundary. In the first case, defects of the scale of the wavelength $\sim\lambda$, breaking such symmetry, are introduced (Fig. 2b). In the second, case a smooth boundary forms a dielectric discontinuity between the disordered medium and free space and, generally, one can expect whispering-gallery (WG) modes due to total internal reflection and cylindrical symmetry [51].
@@
Figure 1: Propagation of a continuous, focused beam through a dense and a sparse layers of irregular particles. (a) and (c), samples with volume fractions of 0.48 (dense) and 0.16 (sparse), respectively. (b) and (d), steady-state near-field intensity distributions $|E|^{2}$ in XY, YZ and back-XZ planes for both layers. The incident beam is $E_{x}$-polarized. The size of particles is $X_{r}=1.1$, and the refractive index is $n=3.0$.
@@
Diffusive propagation of light in sparse media with $l^{*}>\lambda$ is characterized by the transverse spread of energy and an exponential decay of intensity of transmitted light in time as a result of multiple scattering and diffraction by individual scatterers. The geometry of the field spread experiences a change when particles are densely packed. Generally, at dense packing, far-field scattering becomes impossible, and classic multiple scattering turns into percolation of light along random transmission channels [36, 37, 38]. This can lead to formation of Caylee-tree-like channels [36] or a transversely localized propagation [23, 24], depending on the particle size and volume fraction. This transition may occur even at a relatively small refractive index $n=1.5$ and $l^{*}\lesssim\lambda$ [36], which, however, should be insufficient for AL.
@@
In Fig. 1, we study focused beam propagation through thick layers of particles with a size $X_{r}=1.1$ and a refractive index $n=3.0$. The internal field patterns are compared for a dense and a sparse structures of $\rho=0.48$ (Fig. 1a) and 0.16 (Fig. 1c), respectively. A continuous $E_{x}$-polarized focused beam incident on the front side in each case propagates in $Z$ direction. Steady-state intensity distributions as cross-sections in different planes are shown in Figs. 1b and 1d. The ”XY” panels correspond to the cross-sections near the back sides of the samples.
@@
In the sparse case (Fig. 1d), we observe classic scattering by particle groups and individual particles. In a dense system, the size of voids is reduced to the scale of the wavelength or even smaller, and a focused beam tends to preserve its diameter along the path, re-emitting transmitted energy in the center of the back side. Small isotropic spreading occurs due to the leakage of energy through the evanescent field coupling. Therefore, the intensity of the beam decreases, but its geometric cross-section remains constant at the distance of the layer thickness. Excitation of the $E_{y}$ and, particularly, longitudinal $E_{z}$ field components (Fig. S3 in the Supplementary material III) proves that this kind of propagation is not equivalent to the wave transport in a homogeneous medium.
The dynamics of the process can be seen in Video III of the Supplementary material III. It shows the evolution of a short pulse, percolating through a dense layer without transverse spreading. Light propagates mainly through evanescent coupling between individual scatterers without conventional free-space transport of the scattered waves. This reminds us of an evolution of a cellular automata (CA) in which the vector field state in an elementary cell at each time step is determined by its nearest environment, i.e. the shape of neighbour particles or the shape of a sub-wavelength void that is formed by the neighbors. The laws of interaction of the vector field and the air-dielectric interface together with interference become the CA rules. Then, a sequence of such interactions determines the evolution of the entire vector field from the initial state, resulting in a forward pulse propagation.
@@
Figure 2: (a) A layer of 19000 irregular particles packed with a volume fraction of 0.44. The size of particles is $X_{r}=1.1$ and the refractive index is $n=3.0$. (b) vertical cross-section of the layer, showing intentional defects on the side boundary and the characteristic scale of voids $kl\lesssim 1$. (c) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.17$ to $0.5$ as a function of time $T(t)$. (d) Diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. Both axes have logarithmic scale. Dashed line shows a $t^{-1}$ dependence.